A 4:1 molar mixture of He and CH4 is contained in avessel at 20 atm pressure.Due to a hole in vessel, the gas mixture leaks out.What is the composition of the mixture effusing out initially?..........

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
Here,total moles = 4+1 =5,
Total pressure = 20 bars
So, partial pressure of He = 4/5 x 20 = 16 bar Or,
partial pressure of CH = 20-16 = 4 bar
Now, molar mass of He = 4, molar mass of CH = 16 So, ratio of rate of effusion of He (r ) and rate of effusion of methane (r ) will be=8(by calculating)
Hence, as the rate of effusion of He is 8 times higher than CH so, initial mixture must contain the molar ratio of He: CH to be 8:1.